The Tourist Shop sells souvenirs at various locations along the Florida Gulf Coast. During May and June, the company experienced the following events related to t-shirt sales:

May

  1. Purchased inventory for $2,400 on account.
  2. Sold 350 units of inventory for $15 each.
  3. Paid suppliers $3,000.

June

  1. Purchased inventory for $3,600 on account.
  2. Sold 500 units for $15 each.

Related to inventory, The Tourist Shop also recorded the following:

Apr. 30

Ending balance 200 units @ $7

= $1,400

May. 15

Purchased 300 units @ $8

= $2,400

June 5

Purchased 400 units @ $9

= $3,600

The Tourist Shop had the following account balances at April 30:

Accounts payable

$ 2,400

Cash

10,000

Common stock

5,000

Inventory

1,400

Retained earnings

4,000

Required:

  1. Add labels to the first blank line in the horizontal statements model to reflect the components of each statement: Assets, Expenses, Liabilities, Net Income, Revenue, and Stockholder’s Equity.
  2. Add the account titles to the third blank link, grouping them according to the appropriate component of each statement. Add the beginning account balances in the space provided
  3. Record the transactions above in the horizontal statements model provided. The Tourist Shop uses the FIFO method.

What is the total amount of gross profit that The Tourist Shop will report for May and June?
Answer #1
horizontal statements model and monthly gross profit :
Trans. Date Assets = Liabilities + Shareholders' Equity
Cash + Inventory = Accs pay Comm stock Retain. earnings = Revenue-Expenses
Op. Bal. on April 30 10000 1400 = 2400 5000 4000
Purchase May 15 2400 = 2400
Sales May 31 (350*15) 5250 = 5250
COGS May 31 (200*7 + 150*8) -2600 = -2600
Payment May 31 -3000 = -3000                                         
May - Gross Profit = 2650 = 2650
Purchase June 5 3600 = 3600
Sales June 30 (500*15) 7500 = 7500
COGS June 30 (150*8 + 350*9) -4350 = -4350
June - Gross Profit                                                                                   3150 = 3150
Balance 19750 450 = 5400 5000 9800
Grand Balance on June 30 20200 = 20200